Simplify and solve.
15-4(a-5); if a=8
i have tried subsituting "a" with 8, but i cant simplify it that way

Answer by algebrahouse.com(1659)

15 - 4(a - 5), when a = 8
= 15 - 4(8 - 5) {substituted 8 in for a}
= 15 - 4(3) {subtracted 8 - 5}
= 15 - 12 {multiplied 4 by 3}
= 3 {subtracted}

For more help from me, visit: www.algebrahouse.com